how can the number 45 be divided into two parts so that 4 times one part is 9 less than 5 times the other?
Answer by nerdybill(7384)

how can the number 45 be divided into two parts so that 4 times one part is 9 less than 5 times the other?
.
Let x = first part
then
45-x = second part
.
4(45-x) = 5x-9
180 - 4x = 5x-9
180 = 9x-9
189 = 9x
21 = x (first part)
.
Second part:
45-x = 45-21 = 24 (second part)